OpenStudy (anonymous):

Solve for q: mn+q=p (I know how to do this, just making sure I did it right)

terenzreignz (terenzreignz):

Demonstrate, then :)

OpenStudy (anonymous):

Alright, mn+q=p -mn q=p-mn

terenzreignz (terenzreignz):

When solving for a specific variable (in this case, q) Is q alone on one side of the equation? Is there no q on the other side of the equation? If the answer to both questions is yes, then you've done it right :)

OpenStudy (anonymous):

Ok, @terenzreignz thanks :)

terenzreignz (terenzreignz):

No problem.